# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ 2.1 Surgical Workflow Analysis Benchmarks
Over the years, the evolution of datasets for surgical workflow analysis has substantially impacted the field. In 2014, the JIGSAWS dataset (Gao et al. (2014)) emerged as a pioneering effort by combining video and kinematic data with annotations for three fundamental surgical tasks in robotic surgery and fifteen surgical gestures. Nevertheless, this dataset was acquired in a simulated training laboratory, distancing it from an actual surgical setting. Responding to the demand for in vivo human surgical data, the Cholec80 dataset (Twinanda et al. (2016)) introduced a collection of laparoscopic cholecystectomy surgeries with annotations for surgical phase and tool presence recognition. Simultaneously, TUM LapChole (Stauder et al. (2017)), another dataset for laparoscopic cholecystectomies, also labeled surgical phases within a smaller collection of surgeries. Given that the previous datasets focus on traditional laparoscopic procedures, the Nephrec9 dataset (Nakawala (2017)) was introduced as the first surgical workflow analysis dataset with real RAS data of Robot-Assisted Partial Nephrectomy (RAPN) procedures. Remarkably, the Nephrec9 dataset provides temporal annotations at a more granular level, offering information about surgical steps instead of broader surgical phases.
In a different domain, the CATARACTS dataset (Al Hajj et al. (2019)) introduced an annotated video collection tailored for cataract surgeries. The dataset was initially labeled for surgical instrument recognition and expanded to include annotations for surgical phase recognition (Zisimopoulos et al. (2018)) and semantic segmentation of anatomic landmarks and surgical instruments (Grammatikopoulou et al. (2021)), but it does not provide information about surgical actions or steps. Within this same surgical data type, the Cataract-101 dataset (Schoeffmann et al. (2018)) subsequently introduced a collection of 101 videos while still maintaining high-level phase annotations.
A notable advancement in laparoscopic surgery datasets occurred with the introduction of the HeiCo dataset (Maier-Hein et al. (2021)) and the CholecT40, 45, and 50 datasets (Nwoye et al. (2020, 2022); Nwoye and Padoy (2023)). The HeiCo dataset features 30 surgical videos and sensor data from various laparoscopic procedures such as proctocolectomy, rectal resection, and sigmoid resection. The annotations include both phase recognition and surgical instrument segmentation. Remarkably, including diverse surgeries within this dataset allowed for exploring domain generalization in surgical workflow analysis. Nevertheless, it does not include more detailed temporal information, such as action and step recognition, crucial for comprehensive understanding across diverse annotation levels. Contrarily, the CholecT40, 45, and 50 datasets (Nwoye et al. (2020, 2022); Nwoye and Padoy (2023)) used a subset of Cholec80’s data and included annotations for presence recognition of instruments, actions, and anatomical targets via surgical action triplets, thus allowing a more fine-grained study of surgical workflow but still disregarding steps recognition or localized information.
Following this effort, the HeiChole dataset (Wagner et al. (2023)) presented a detailed collection of 33 laparoscopic cholecystectomies and is also aimed at multi-level understanding by presenting phase recognition along with frame-level instrument and action presence recognition. While more specialized in its scope than the HeiCo dataset, HeiChole did not include crucial spatially localized information or step recognition within surgical procedures.
Even though the previously mentioned datasets have offered valuable insights into surgical workflow analysis within their specific surgical domains, they all confine their video-level annotations to studying a single long-term task. The MISAW dataset (Huaulmé et al. (2021)) emerged as a crucial addition to surgical workflow analysis to address the previous gap in existing datasets. Focused on micro-surgical anastomosis on artificial blood vessels, MISAW encompassed 27 sequences with videos, kinematic data, and temporal annotations across three granularity levels: phases, steps, and specific activities that are constituted by a surgical gesture, a target, and a surgical instrument. This dataset marked a significant stride towards enabling multi-level surgical workflow analysis. However, the artificial nature of the data limits its applicability to real-time surgical scene understanding.
In contrast, the RARP-45 dataset (van Amsterdam et al. (2022)) stands out as a significant contribution by addressing the need for finer-grained temporal annotation within human in vivo robot-assisted interventions. This dataset comprises a comprehensive collection of video and kinematics data derived from Robot-Assisted Radical Prostatectomies (RARP), enriched with temporal segment annotations for fine-grained bi-manual gesture recognition. Nevertheless, it solely offers surgical gesture annotations. Within the same RAS domain, the recently introduced RARR dataset (Kolbinger et al. (2023)) allows multi-level understanding by introducing annotations for higher-level phase recognition and lower-level anatomical targets segmentation in human in vivo robot-assisted rectal resection (RARR). Despite providing spatial and temporal annotations, it does not address instrument recognition or the multi-granularity of broken-down phases into steps or surgical actions.
In exploring surgical workflow analysis advancements, we acknowledged the need for a single benchmark encompassing hierarchically various granularity levels of long- and short-term surgical workflow tasks. We especially recognize this need in datasets utilizing real RAS data on human subjects. Our dataset, as an extension of our previously proposed PSI-AVA dataset (Valderrama et al. (2022)), provides annotations with high-level temporal information regarding phases and their division into steps, a more detailed temporal understanding. Thus, ours is the only surgical workflow analysis dataset up to date that provides both levels of semantic understanding with human in vivo data. Additionally, our dataset contains detailed low-level spatio-temporal information on atomic actions, and, in this work, we include precise delineation of the different surgical instruments for a more detailed understanding of instrument poses and surgical gestures. Consequently, our work addresses multiple previous needs in surgical scene understanding by offering a holistic comprehension of the surgical workflow at different granularity levels with real RAS data on human subjects.
2.2 Surgical Instrument Segmentation
Most benchmarks for surgical instrument segmentation started as part of the Endoscopic Vision Grand Challenge [^4] https://endovis.grand-challenge.org/. The inaugural benchmark in this series was the 2015 Instrument Segmentation and Tracking Challenge (Endovis 2015) (Bodenstedt et al. (2018)), which provided short video clips of laparoscopic colorectal surgeries and some ex vivo robot-assisted procedures. The corresponding semantic segmentation annotations focused on the sub-parts of the surgical instruments. Later, the 2017 Robotic Instrument Segmentation Challenge (Endovis 2017) (Allan et al. (2019)) introduced a more complex approach to instrument segmentation by proposing the recognition of instrument types. Thus, Endovis 2017 provided instance-wise annotations for entire instruments and their sub-type label in multiple videos of robot-assisted porcine procedures. The 2018 Robotic Scene Segmentation Challenge (Endovis 2018) (Allan et al. (2020)) further expanded the scope of these benchmarks, offering surgical scene semantic segmentation in robot-assisted porcine procedures. These detailed annotations included instrument parts, anatomical structures, and specific surgical objects like clips, threads, and needles, which provided a richer context for segmentation tasks. Still, all these benchmarks solely focus on segmentation tasks, lacking annotations for surgical workflow analysis tasks, and none of them provide data from human in vivo robot-assisted procedures. Also, all these benchmarks were mainly formulated for semantic segmentation approaches, as they all use semantic segmentation-based annotations and metrics, and only Endovis 2017 could be directly employed for instance segmentation. Thus, these benchmarks significantly ignored the instance-level nature of surgical instrument segmentation.
In our previous research (González et al. (2020)), our team adapted the original annotations from Endovis 2018 for instance-level segmentation of instrument types. We used these annotations to fully exploit the potential of mask classification models for instance segmentation (He et al. (2017)). This approach demonstrated the superior potential of instance-based approaches for surgical instrument segmentation, as we strongly outperformed all previous semantic segmentation-based methods (González et al. (2020)). Nevertheless, we kept a semantic segmentation evaluation metric to maintain comparability with previous studies. Alternatively, Islam et al. (2020) re-annotated the data from Endovis 2018 for detecting instruments and anatomical structures, along with annotations for the interactions between instruments and tissues. These interaction annotations are analogous to instrument action annotations and provide a deeper understanding of the dynamics within the surgical scene.
Finally, the CaDIS dataset (Grammatikopoulou et al. (2021)) and the HeiCo dataset (Maier-Hein et al. (2021)) were the first ones to provide instrument segmentation and video-level task annotations. The CaDIS dataset is a subset of the CATARACTS dataset (Al Hajj et al. (2019)), which focuses on cataract surgery and provides phase recognition annotations. CaDIS offers entire scene semantic segmentation that includes instrument types and anatomical structures. Further, the 2019 ROBUST-MIS challenge (Roß et al. (2021)) used the videos of the HeiCo dataset (Maier-Hein et al. (2021)) (also with phases recognition annotations) for instance-level instrument-type segmentation and established the first instance-based evaluation metric. Still, these datasets use different data sources from robot-assisted surgery and do not cover relevant recognition tasks like steps or action recognition. Our dataset aims to bridge the gaps mentioned above, offering pixel-level annotations for robotic surgical instrument segmentation, along with the atomic actions performed by those instruments and multiple video-level recognition tasks in real robot-assisted surgical footage.
@@ 2.3 Surgical Action Recognition
The concept of action recognition in computer vision began as a classification task in short clips (Schuldt et al. (2004); Blank et al. (2005); Marszalek et al. (2009); Kuehne et al. (2011); Soomro et al. (2012) and extended videos (Over et al. (2013); Karpathy et al. (2014); Abu-El-Haija et al. (2016); Kay et al. (2017); Goyal et al. (2017); Zhao et al. (2019)), where a single action category was associated with a specific time segment. Over time, the problem expanded to multiple categories within each time segment (Caba Heilbron et al. (2015); Idrees et al. (2017); Yeung et al. (2018); Sigurdsson et al. (2016)). Some works further extended the problem to include spatial location detection of actions (Ke et al. (2005); Rodriguez et al. (2008); Yuan et al. (2009); Jhuang et al. (2013)). Subsequently, AVA (Gu et al. (2018)) introduced an agent-centric approach, suggesting that a video consisted of various agents capable of independently executing multiple actions and interacting with each other and their environment. This approach incorporated atomic actions, which divided categories into finer components for a richer and more complete understanding of actions in videos.
In the surgical context, action recognition corresponds to a higher temporal granularity in surgical scene understanding (Khatibi (2020); Rupprecht et al. (2016)). The JIGSAWS dataset (Gao et al. (2014)) made early efforts in this direction by defining actions as gestures within the surgical environment. This dataset adopted a classification framework, including 15 distinct categories observed in ex vivo surgery videos recorded by the da Vinci Surgical System (dVSS). However, as mentioned earlier, it derives its data from external sources to the patient, limiting its representativeness in real-life cases
Subsequently, Sharghi et al. (2020) introduced a video dataset captured within an operating room, featuring ten specific actions performed by medical personnel during robot-assisted surgeries. Nevertheless, this dataset does not include scenarios involving in-patient data. Later, the CholecT40 dataset (Nwoye et al. (2020)) and its subsequent versions, CholecT45 and CholecT50 (Nwoye et al. (2022); Nwoye and Padoy (2023)), formulated the challenge of action identification in laparoscopic cholecystectomy surgeries as triplets. The task was defined as an instrument that executes an action on a target, adopting the structure of $<$instrument, verb, target$>$. Correspondingly, the MISAW dataset (Huaulmé et al. (2021)) employed a similar formulation for anastomosis micro-surgeries. However, the use of triplets does not explicitly provide the location where each action takes place, which would allow for a more detailed understanding of surgical scenes.
Recently, the RARP45 (van Amsterdam et al. (2022)) and the HeiChole (Maier-Hein et al. (2021)) datasets introduced action recognition frameworks, approaching them as frame-wise recognition problems similar to that of JIGSAWS, but within the context of in vivo scenarios. RARP45 includes ten distinct actions on videos of Robot-Assisted Radical Prostatectomies. Similarly, HeiChole extends this approach to laparoscopic cholecystectomy with four possible categories. However, these annotations do not include spatial location information and offer a rather generalized understanding of actions.
@@
Furthermore, Islam et al. (2020) introduced the localization of actions within surgical scenes in the Endovis 2018 dataset to enhance comprehension of the implicit relationships between instruments and tissues. The approach associated interaction categories with bounding boxes in Endovis 2018, designating the specific actions executed by each instrument or tissue. The dataset annotations used a graph structure, where each node represented an instrument or tissue, and the edges depicted the interactions between these nodes. Subsequently, the ESAD dataset (Bawa et al. (2021)) extended the localization of actions in robot-assisted surgery to Radical Prostatectomy data. Moreover, the AVOS dataset (Goodman et al. (2021)) expanded action localization to multiple surgery types. Nonetheless, these datasets assign a single action category to each instrument and offer classes that could be further analyzed into even finer actions.
@@
Contrary to previous approaches, our work adopts the Atomic Visual Actions (AVA) (Gu et al. (2018)) formulation for a more detailed study of temporal activities by employing atomic actions. As their name suggests, atomic actions are indivisible and represent the most fine-grained level of action understanding, thus allowing practical and precise activity modeling. Our GraSP dataset and its predecessor are pioneering in integrating the concept of atomic actions into the surgical domain. Accordingly, our benchmark deconstructs surgical activities into various atomic actions that surgical tools perform. As a result, GraSP presents a localized and multi-target atomic action detection task where each instrument independently performs multiple atomic actions simultaneously. This novel and highly challenging approach facilitates a more in-depth analysis of surgical actions and endorses applications across diverse surgery types, as it is a naturally transferable notion to any surgical domain.
@@ 2.4 Surgical Scene Understanding with Vision Transformers
Transformer models (Vaswani et al. (2017)) made a breakthrough in natural language processing, demonstrating their formidable capacity to process sequences by employing attention mechanisms across temporal dimensions. With the introduction of Vision Transformers (Dosovitskiy et al. (2021)), their application expanded to the realm of visual data processing. In recent years, Vision Transformers have gained significant prominence, showcasing their ability to utilize attention mechanisms to capture intricate and wide-ranging dependencies among diverse elements within an image. This ability translated to encouraging outcomes in image and video analysis (Carion et al. (2020); Fan et al. (2021); Zhu et al. (2021)).
This success has extended to the domain of endoscopic surgeries since previous works have successfully included transformers in their approach. For instance, in the domain of instrument classification, Kondo (2021) proposed a technique employing a CNN for feature extraction along with a temporal module based on Transformers, offering insights into a temporal window of features. In a distinctive approach, Nwoye et al. (2022) introduced a Transformer-inspired method for classifying triplets, incorporating a Convolutional Neural Network (CNN) backbone and attention mechanisms for enhanced feature mapping. Later, Sharma et al. (2023a) expanded on this model by incorporating a temporal consistency module, which fused features within a temporal window. Recently, Sharma et al. (2023b) proposed a triplet detection model, leveraging Deformable DETR (Zhu et al. (2021)) and a ResNet50 (He et al. (2016)) coupled with a transformer encoder for feature extraction. The methodology included incorporating a graph to process these features and understand interactions and a Multi-Head Attention layer applied to the nodes’ features.
In phase recognition, Gao et al. (2021) utilized a CNN backbone for spatial feature extraction, coupled with temporal processing through a Temporal Convolutional Network (TCN) and refinement with Transformer layers. Similarly, Czempiel et al. (2021) integrated a CNN backbone and Transformer layers, introducing a regularization parameter to focus attention on high-confidence features. Additionally, Ding and Li (2022) explored phases as temporal segments, employing a combination of CNN, temporal convolutions, and Transformers to achieve an enhanced capture of semantics and interrelations.
Addressing action localization, Zhang et al. (2022) compared models based on CNN, Transformers, and a hybrid approach. Superior results were achieved by integrating a CNN-based backbone with a BERT-based transformer network (Devlin et al. (2019)), showcasing enhanced temporal video understanding compared to ViT (Dosovitskiy et al. (2021)) for feature extraction.
Regarding the task of instrument segmentation, Zhao et al. (2022) pioneered a region classification approach, utilizing a modified MaskFormer (Cheng et al. (2021)) along with a CNN backbone. However, this model is limited in its comprehension of long-term surgical dynamics, relying solely on the previous frame for predictions in the current frame. Recently, Dhanakshirur et al. (2023) employed a Mask DINO (Li et al. (2023)) model with a ResNet50 (He et al. (2016)) backbone for instrument segmentation. The approach incorporated a transformer encoder to process the backbone features and implemented layers of various attention forms using the best queries predicted by Mask DINO to enhance the consistency and quality of region proposals.
However, these methodologies have not exploited the potential of transformers for a comprehensive understanding of surgical scenes, often relying on CNNs for feature extraction. Moreover, these approaches are designed for specific tasks rather than employing a general design that embraces the holistic nature of surgeries. Regarding these gaps, our previous works, Valderrama et al. (2022) and Ayobi et al. (2023), leveraged transformers for various tasks. These methods marked the introduction of architectures entirely based on Transformers, demonstrating satisfactory results in instrument detection, instrument segmentation, phase classification, and step classification tasks, achieving state-of-the-art performance. In this regard, our TAPIS model introduces a single method fully based on transformer architectures that can effectively tackle multiple surgical workflow analysis tasks within different levels of visual, temporal, and spatial understanding. Thus presenting a generalized state-of-the-art transformer model for multi-granular surgical scene understanding.

### v3

@@ 2.1 Surgical Workflow Analysis Benchmarks
Table 1 summarizes key SWA benchmarks introduced over the years along with their respective semantic tasks. Early efforts primarily focused on temporal analysis with independent video or frame-level tasks. The JIGSAWS dataset ([29]) used videos from simulated robot surgeries annotated for surgical skill assessment and gesture recognition. The Cholec80 dataset ([83]) first introduced in vivo human data of laparoscopic cholecystectomies with annotations for surgical phase and tool recognition. Simultaneously, M2CAI16 ([79]) used similar data with surgical phase labels only. The Nephrec9 dataset ([62]) first introduced human Robot-Assisted Partial Nephrectomy data with finer labels for surgical step recognition rather than phases. However, none of these datasets provided any spatial annotations.
Furthermore, the HeiCo ([60]) and AutoLaparo ([89]) datasets feature various laparoscopic procedures with temporal and spatial annotations for phase recognition and surgical instrument segmentation, but lack more granular temporal tasks. The CholecT50 dataset ([63, 65, 64]) used part of Cholec80’s data with additional test videos and annotations for presence recognition of instruments, actions, and anatomical targets using surgical action triplets. Similarly, the HeiChole dataset ([87]) encompasses laparoscopic cholecystectomies and presents annotations for phases, actions, and instrument recognition. While CholecT50 and HeiCHole allow for finer-grained SWA, they still disregard spatially localized information or step recognition.
Recent frameworks have partially addressed previous gaps in multi-granularity modeling and the availability of RAS data. The MISAW dataset ([36]) focuses on artificial micro-surgical anastomosis, offering temporal annotations across three granularity levels: phases, steps, and specific activities, though it lacks spatial information. In contrast, the RARP-45 dataset ([4]) comprises videos from human Robot-Assisted Radical Prostatectomies with temporal annotations for gesture recognition, and spatial annotation for instrument segmentation ([68]). The RARR dataset ([47]) introduces annotations for phase recognition and anatomical target segmentation in human in vivo robot-assisted rectal resection. However, RARP45 and RARR only account for a single temporal task.
Conversely, our dataset uniquely combines annotations for high-level surgical phases and their decomposition into steps, making it the only SWA dataset to offer both levels of temporal semantic understanding using human in vivo data. Additionally, it incorporates detailed annotations for atomic actions and instrument segmentation, thus addressing previous gaps in temporal and spatial multi-granularity and presenting a more holistic approach to SWA.
2.2 Surgical Instrument Segmentation Benchmarks
Most benchmarks for surgical instrument segmentation originated from the Endoscopic Vision Grand Challenge [^4] https://endovis.grand-challenge.org/. First, the 2015 Instrument Segmentation and Tracking Challenge (Endovis 2015) ([9]), offered laparoscopic colorectal surgeries and ex vivo robot-assisted procedures with semantic segmentation of instrument sub-parts. Later, the 2017 Robotic Instrument Segmentation Challenge (Endovis 2017) ([3]) introduced instrument type recognition with instance-wise annotations in robot-assisted porcine procedures. Similarly, The 2018 Robotic Scene Segmentation Challenge (Endovis 2018) ([2]) used similar data with semantic segmentation annotations for instrument parts, anatomical structures, and additional surgical objects. Nevertheless, these benchmarks lack higher-level temporal tasks, do not include human in vivo RAS data, and, in the cases of EndoVis 2015 and 2018, they overlook the instance-level nature of surgical instrument segmentation.
In our previous research ([32]), we adapted Endovis 2018’s annotations for instance-level segmentation of instrument types and demonstrated the superior effectiveness of instance-based approaches for surgical instrument segmentation. Further, the 2019 ROBUST-MIS challenge ([72]) used HeiCo’s ([60]) data to establish the first instance-based evaluation for instrument-type segmentation, which was subsequently employed by AutoLaparo ([89]) on laparoscopic hysterectomies. More recently, the SAR-RARP50 ([68]) challenge expanded RARP-45’s ([4]) data and introduced the first segmentation dataset with human RAS data for semantic segmentation of instrument parts and surgical objects. Consequently, no prior dataset offers instance-level instrument type segmentation on human RAS data, making our benchmark the first to provide this critical task.
@@ 2.3 Surgical Action Recognition
In the surgical context, action recognition represents a finer temporal task than phases or steps ([45, 73]), and several formulations have been developed to model this task. The JIGSAWS dataset ([29]) first defined surgical actions as gestures, which represent short temporal segments of discrete surgical activity units, an approach later adopted by the RARP-45 ([4]) benchmark. Subsequently, the CholecT50 dataset ([63, 65]) formulated surgical actions as multiple frame-wise action triplets in the structure of $<$instrument, verb, target$>$, which were also utilized by the MISAW dataset ([36]). The HeiChole ([60]) datasets approached action recognition as a frame-wise presence recognition problem on a small set of atomic action categories. However, these benchmarks do not provide the explicit location of actions, which would allow for a more detailed understanding of surgical scenes.
@@
Furthermore, [39] introduced action localization within surgical scenes in the Endovis 2018 dataset using a graph structure, where nodes represented instrument or tissue bounding boxes, and edges depicted the interactions between instruments and tissues. Subsequently, the ESAD dataset ([8]) extended the localization of actions to Robot-Assisted Radical Prostatectomy data, using bounding boxes around visible instruments and labeling their primary actions. Nonetheless, these datasets assign each instrument a single action category and offer classes that could be further analyzed to identify even finer actions.
@@
Contrary to previous approaches, our work adopts the Atomic Visual Actions ([34]) formulation for a more detailed study of temporal activities. By definition, atomic actions are indivisible, agent-centered, and represent the finest level of action granularity. Accordingly, our benchmark deconstructs surgical activities into various independent atomic actions that each surgical tool performs simultaneously. As a result, GraSP presents a localized and multi-target atomic action detection task, a novel and highly challenging approach that facilitates more in-depth analysis of surgical actions and is a naturally transferable notion to any surgical domain.
@@ 2.4 Surgical Scene Understanding with Vision Transformers
In recent years, transformer models ([85, 20]) have gained prominence for their ability to capture long-range dependencies using attention mechanisms. These models offer a superior modeling capabilities that consistently surpass Convolutional Neural Networks (CNNs) in image and video analysis ([20, 11, 24, 94]). This success has impacted the domain of endoscopic surgeries. For instance, LapFormer ([48]) and Opera ([16]) employed CNN-based frame feature extractors with temporal consistency transformers for instrument and phase recognition, respectively. Moreover, Trans-SVNet ([28]) and SAHC ([19]) used transformer layers to refine the outputs of a CNN backbone and a Temporal Convolutional Network for phase recognition. Recently, LoVit and SKiT ([53, 54]) used a fully transformer-based, two-stage architecture for phase recognition, leveraging a Vision Transformer (ViT) ([20]) backbone and a long-term transformer. However, these phase recognition methods rely on global frame-wise encoders to produce generalized frame embeddings for long-range temporal architectures, hence limiting their adaptability for more granular tasks that demand finer spatial and temporal embeddings.
Furthermore, transformers have been utilized to improve surgical action triplet recognition. Rendezvous ([65]) employs a CNN backbone with an adapted transformer encoder, while Rendezvous in Time (RiT) ([76]) further enhances the encoder by integrating a temporal consistency module. In contrast, MCIT-IG ([77]) leveraged a transformer detector ([94]) coupled with a CNN backbone and a modified transformed decoder to improve triplet identification. Nevertheless, these models also utilize frame-wise encoders and have limited temporal processing capabilities. While these models can process certain levels of activity localization ([65]), they do not explicitly output instruments or action regions, which limits their effectiveness for detailed spatial analysis.
Regarding instrument segmentation, TraSeTR ([92]) utilized a modified MaskFormer ([13]) with a CNN backbone, integrating instrument region embeddings of previous frames to improve temporal modeling. Furthermore, [18] employed a Mask DINO ([51]) model, combining a CNN backbone with a Query Proposal Network to enhance region proposals. Recent approaches, such as [93, 90], employ prompting strategies for foundation transformer models to improve generalizability, but these methods differ from the supervision approach of this work. Still, current segmentation architectures use very limited temporal context and cannot adapt to localized temporal tasks like atomic actions.
Therefore, previous methodologies are tailored to specific surgical scene understanding tasks, often relying on frame-wise CNN feature extractors. Consequently, no fully transformer-based architecture has been developed to adapt to different tasks while capturing the holistic and multi-granular nature of surgeries, therefore limiting the exploitation of transformers’ powerful modeling capabilities. Regarding these gaps, our previous works, [84] and [5], demonstrated outstanding results on diverse tasks by using a fully transformer-based model. In this regard, our TAPIS model introduces a transformer method incorporating a video-wise feature encoder and a region proposal network. This design enables TAPIS to effectively handle multiple surgical workflow analysis tasks at various levels of visual, temporal, and spatial understanding. Thus, it presents the first generalized state-of-the-art transformer model for understanding multi-granular surgical scenes.